A firm that owns the stock of another corporation does not have to pay taxes on the entire amount of dividends received. In general, only 30 percent of the dividends received by one corporation from another are taxable. The purpose of this tax law feature is to mitigate the effect of triple taxation, which occurs when earnings are first taxed at the first firm, its dividends paid to the second firm are taxed again, and the dividends paid to shareholders by the second firm are taxed yet again. Assume that a firm with a 21 percent tax rate receives $100,000 in dividends from another corporation. What taxes must be paid on this dividend, and what is the after-tax amount of the dividend?
Answer #1
Dividend received= $100,000

Only 30% are taxable

So taxable portion of dividend = $100000*30%= $30,000
Tax rate= 21%
So tax on dividends = 30000*21%= $6,300

Net dividend after tax = Total dividend - tax paid

100000-6300
$93,700

So taxes must be paid i $6300

After tax amount of dividend is $93700

Please thumbs up